OpenStudy (hungergames13):

A line has a slope of −5 and contains the point (−2, 4). Which equations represent the line? Choose all answers that are correct. A. −5x + y = −6 B. 5x + y = −6 C. (y – 4) = −5(x + 2) D. (y + 4) = −5(x − 2)

OpenStudy (welshfella):

the point slope form of the equation of a straight line is y - y1 = m(x - x1) where the slope = m and (x1,y1) is a point on the line

OpenStudy (welshfella):

so if you plug in the values you have into the above you''l get the required eqution

OpenStudy (welshfella):

- well one of them

OpenStudy (hungergames13):

ohh so it's D?

OpenStudy (welshfella):

No

OpenStudy (welshfella):

y1 = 4 so the left side is y - 4

OpenStudy (anonymous):

it would be c and b because you can get those by the equation that welshfella showed. then the other way you get it by graphing. it would not be d because you plug 4 into the equation there fore it would be y-4=-5(x+2)

OpenStudy (welshfella):

you can also get b by rearranging equation c.

OpenStudy (hungergames13):

ohhh, ok I get it, thank you soo much.

OpenStudy (welshfella):

yw